OpenStudy (anonymous):

I WILL GIVE A MEDAL!! A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2 and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold.

OpenStudy (anonymous):

1.Change the equation into slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all of your work. 2.Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences. 3.Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences. 4.Graph the function using one of the following two options below. One the graph, make sure to label the intercepts. ~You may graph your equation by hand on a piece of paper and scan your work. ~You may graph your equation using graphic technology that can be found in the Course Information area. 5.Suppose in the next month, the total profit on smart phone cases and tablet cases is $1,500. The profit amounts are the same, $2 for smart phone case and $3 for the tablet case. In a paragraph of at least three sentences, explain how the graphs of the functions for the two months are the same and how they are different. Be sure to use complete sentences. 6.Below is a graph that represents the total profits for a third month. Write the equation of the line that represents this graph. Show your work or explain how you determined the equations.

OpenStudy (anonymous):

1. \[y = \left(\begin{matrix}2 \\ 3\end{matrix}\right)x+400\]

OpenStudy (anonymous):

@EEmarr could you help me with the rest of them?

undeadknight26 (undeadknight26):

I did this same assignment let me tell you what i had and how i got there...

OpenStudy (anonymous):

i had to redo this

undeadknight26 (undeadknight26):

ahh me too ok one second...what teacher?

OpenStudy (anonymous):

Larken

undeadknight26 (undeadknight26):

Oh cool i got Lohr.

OpenStudy (anonymous):

cool

undeadknight26 (undeadknight26):

You in OCVS?

OpenStudy (anonymous):

flvs

undeadknight26 (undeadknight26):

1. 2x + 3y = 1200 2x-1200=-3y 1/3*(2x-1200)=1/3*-3y 2/3x-400=-y -2/3x+400=y

undeadknight26 (undeadknight26):

The second one you would need to find the slope intercept which was given above, -2/3... 2. I would put a dot on (0,400) then I would use the slope to do the rest or go 2 points to the right and 3 to the down. After this it should look like 3,398.

undeadknight26 (undeadknight26):

Do not use my exact wording...try re wording it or saying if differently.

OpenStudy (anonymous):

ok

undeadknight26 (undeadknight26):

3. F(x) = 2x + 1200

undeadknight26 (undeadknight26):

Can you do the graph by yourself?

undeadknight26 (undeadknight26):

Sorry i have it on paper so i would have to scan it again and stuff...

undeadknight26 (undeadknight26):

5. 2x + 3y = 1500 -3y = 2x – 1500 1/3* -3y = (2x – 1500)*1/3 -2/3x + 500 = y

undeadknight26 (undeadknight26):

6. F(x)=mx + b F(x) = mx + 300 F(x) = 6/9x + 300 F(x) = 2/3x + 300

undeadknight26 (undeadknight26):

That should be it....

OpenStudy (anonymous):

thx this helped me out alot

undeadknight26 (undeadknight26):

Please review this and learn :D and a pleasure to help you.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

This helped me to, Thanks :)

undeadknight26 (undeadknight26):

No problem person XD

OpenStudy (stormswan):

This ALSO helped me. you're awesome @undeadknight26

OpenStudy (anonymous):

what about 4 @johnweldon1993

OpenStudy (anonymous):

@djcool31 @christiansousley

OpenStudy (anonymous):

thanks im doing it right now but i need help with 4 how do i graph @undeadknight26

undeadknight26 (undeadknight26):

Try desmos graphing or make a graph and plot it! I used desmos because it was easier :D Just enter the equation and it will graph it for you!

OpenStudy (anonymous):

ok thanks @undeadknight26

undeadknight26 (undeadknight26):

Anytime! Welcome to the site by the way! Have fun, be safe, be nice!

OpenStudy (anonymous):

thanks

OpenStudy (secret-ninja):

Thanks @undeadknight26 ! You're a lifesaver as usual! :D

undeadknight26 (undeadknight26):

Anything for you miss!

OpenStudy (confluxepic):

I searched a different question. Found this.

undeadknight26 (undeadknight26):

Well than you musta been surprised then mate XD

OpenStudy (confluxepic):

I was.

OpenStudy (secret-ninja):

*watches*

OpenStudy (secret-ninja):

Bye.

OpenStudy (confluxepic):

Bye.

undeadknight26 (undeadknight26):

Welp. That was fun and not intense at all ._.

OpenStudy (confluxepic):

I know right.

OpenStudy (shes.radical):

You . Rock . You dont even know how much you just helped me .. 1 year later cx @undeadknight26 Thanks !!

undeadknight26 (undeadknight26):

Any time @shes.radical <3

OpenStudy (shes.radical):

Heeey .. @undeadknight26 For #3 , It says to explain what the graph of the function represents .. What does the graph of the function represent ? Lol , sorry but I dont get it .

OpenStudy (darronw):

@Secret-ninja @shes.radical @confluxepic Are you guys still in flvs too? what chapter are you in? We can message on there if we're in the same class, does anyone have ms prince?

OpenStudy (confluxepic):

@darronw I am not in FLVS.

OpenStudy (anonymous):

i need help with number 4 cna you help me plz? @undeadknight26

OpenStudy (darronw):

@Calvin3 who's your teacher? as long as we don't have the same one I can help

undeadknight26 (undeadknight26):

For number 4 all you need to do is make the graph :D

undeadknight26 (undeadknight26):

Use: Desmos.com/calculator It's very useful!